OpenStudy (anonymous):

find all solutions to the equations 2sin(theta)+1=0 write answer in terms of pie =)

OpenStudy (angela210793):

2sinx=-1 sinx=-1/2 x=2n*pi-pi/2

OpenStudy (angela210793):

x=theta

OpenStudy (anonymous):

whats n?

OpenStudy (anonymous):

N is any integer 0,1,2,3... There are infinitely many solutions since the trig functions are periodic.

OpenStudy (anonymous):

2sin(theta)+1=0 sin(theta)=-1/2 theta=-pi/6

OpenStudy (angela210793):

yea,thnx Malevolence :)

OpenStudy (anonymous):

-pi/6=30

OpenStudy (anonymous):